OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}+4y_p^\prime+4y_p=12x^2e^{-2x}\]

OpenStudy (unklerhaukus):

\[y_p=(C+Dx)xe^{-2x}\] \[y^\prime_p=(Dx+C+Dx-2x(C+Dx))e^{-2x}\]\[\qquad\quad =(Dx+C+Dx-2Cx-2Dx^2)e^{-2x}\]\[\qquad\quad =(C+2Dx-2Cx-2Dx^2)e^{-2x}\] \[y^{\prime\prime}_p=\left(2D-2C-4Dx-2(C+2Dx-2Cx-2Dx^2)\right)e^{-2x}\]\[\qquad\qquad\quad=\left(2D-2C-4Dx-2Dx-2C-2Dx+4Cx+4Dx^2\right)e^{-2x}\]\[\qquad\qquad\quad=\left(2D-4C-8Dx+4Cx+4Dx^2\right)e^{-2x}\] \[\begin{array}{ccc}\left(2D-4C-8Dx+4Cx+4Dx^2\right)e^{-2x}\\+4(C+2Dx-2Cx-2Dx^2)e^{-2x}\\+4(C+Dx)xe^{-2x}\end{array}=12x^2e^{-2x}\] \[\left(\begin{array}{ccc} 2D-4C-8Dx+4Cx+4Dx^2 \\+4C+8Dx-8Cx-8Dx^2\\ \qquad\qquad+4Cx+4Dx^2\end{array}\right)e^{-2x}=12x^2e^{-2x}\] \[2D=12x^2\]

OpenStudy (unklerhaukus):

something went wrong

OpenStudy (lalaly):

\[y_p \text{ should be } y_p=Cx^2e^{-2x}+Dx^3e^{-2x}+Ex^4e^{-2x}\]since e^-2x and xe^(-2x) is in the homogeneous solution

OpenStudy (unklerhaukus):

agh

OpenStudy (unklerhaukus):

?

OpenStudy (anonymous):

\[y_p=Cx^2e^{-2x}+Dx^3e^{-2x}+Ex^4e^{-2x}\] \[y_p'=2Cxe^{-2x}-2Cx^2e^{-2x}+3Dx^2e^{-2x}-2Dx^3e^{-2x}+4Ex^3e^{-2x}-2Ex^4e^{-2x}\] \[y_p''=2Ce^{-2x}-4Cxe^{-2x}-4Cxe^{-2x}+4Cx^2e^{-2x}+6Dxe^{-2x}-6Dx^2e^{-2x}-6Dx^2e^{-2x}\]\[+4Dx^3e^{-2x}+12Ex^2e^{-2x}-8Ex^3e^{-2x}-8Ex^3e^{-2x}+4Ex^4e^{-2x}=\] \[2Ce^{-2x}-8Cxe^{-2x}+4Cx^2e^{-2x}+6Dxe^{-2x}-12Dx^2e^{-2x}+4Dx^3e^{-2x}\]\[+12Ex^2e^{-2x}-16Ex^3e^{-2x}+4Ex^4e^{-2x}\]

OpenStudy (unklerhaukus):

isn't this a better particular solution? \[y_p=(Cx+Dx^2+Ex^3)e^{-2x}\]

OpenStudy (unklerhaukus):

i guess not, dammnit

OpenStudy (unklerhaukus):

\[y_p=(Cx^2+Dx^3+Ex^4)e^{-2x}\] \[y^\prime_p=\left(2Cx+3Dx^2+4Ex^3-2(Cx^2+Dx^3+Ex^4)\right)e^{-2x}\]\[\qquad\quad =\left(2Cx+3Dx^2+4Ex^3-2Cx^2-2Dx^3-2Ex^4\right)e^{-2x}\]

OpenStudy (unklerhaukus):

\[\Tiny y^{\prime\prime}_p=\left(2C+6Dx+12Ex^2-2Cx-6Dx^2-8Ex^3-2(2Cx+3Dx^2+4Ex^3-2Cx^2-2Dx^3-2Ex^4)\right)e^{-2x}\]\[\Tiny \qquad=\left(2C+6Dx+12Ex^2-2Cx-6Dx^2-8Ex^3-4Cx-6Dx^2-8Ex^3+4Cx^2+4Dx^3+4Ex^4\right)e^{-2x}\]\[\Tiny \qquad=\left(2C+6Dx+12Ex^2-6Cx-12Dx^2-16Ex^3+4Cx^2+4Dx^3+4Ex^4\right)e^{-2x}\]

OpenStudy (unklerhaukus):

our derivatives are in agreement

OpenStudy (anonymous):

yours has -6Cx while mine is -8Cx *checking*

OpenStudy (unklerhaukus):

oh yeah

OpenStudy (unklerhaukus):

-8Cx, will cancel better,

OpenStudy (unklerhaukus):

i see my mistake , it on the first line of the second derivative the fourth term -2Cx should be-4Cx

OpenStudy (unklerhaukus):

\[\tiny y^{\prime\prime}_p=\left(2C+6Dx+12Ex^2-4Cx-6Dx^2-8Ex^3-2(2Cx+3Dx^2+4Ex^3-2Cx^2-2Dx^3-2Ex^4)\right)e^{-2x}\]\[\tiny\qquad=\left(2C+6Dx+12Ex^2-4Cx-6Dx^2-8Ex^3-4Cx-6Dx^2-8Ex^3+4Cx^2+4Dx^3+4Ex^4\right)e^{-2x}\]\[\tiny\qquad=\left(2C+6Dx+12Ex^2-8Cx-12Dx^2-16Ex^3+4Cx^2+4Dx^3+4Ex^4\right)e^{-2x}\]

OpenStudy (anonymous):

yea

OpenStudy (unklerhaukus):

\[\begin{array}{ccc}\left(2C+6Dx+12Ex^2-8Cx-12Dx^2-16Ex^3+4Cx^2+4Dx^3+4Ex^4\right)e^{-2x} \\+4\left(2Cx+3Dx^2+4Ex^3-2Cx^2-2Dx^3-2Ex^4\right)e^{-2x} \\+4(Cx^2+Dx^3+Ex^4)e^{-2x}\end{array}=12x^2e^{-2x}\] \[\left(\begin{array}{ccc} 2C+6Dx+12Ex^2-8Cx-12Dx^2-16Ex^3+4Cx^2+4Dx^3+4Ex^4 \\ \qquad\qquad\qquad\qquad+8Cx+12Dx^2+16Ex^3-8Cx^2-8Dx^3-8Ex^4 \\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+4Cx^2+4Dx^3+4Ex^4\end{array}\right)e^{-2x}=12x^2e^{-2x}\] \[(2C+6Dx+12Ex^2)=12x^2\]

OpenStudy (unklerhaukus):

C=0 D=0 E=1

OpenStudy (unklerhaukus):

\[y_p=x^4e^{-2x}\]

OpenStudy (unklerhaukus):

\[y(x)=(A+Bx)e^{-2x}+x^4e^{-2x}\qquad\checkmark\]

OpenStudy (unklerhaukus):

thanks for your help @nphuongsun93

OpenStudy (anonymous):

ty ty